Lead time for one of your fastest- moving products is 21 days. Demand during this period averages 100 units per day.

a) What would be an appropriate reorder point?

b) How does your answer change if demand during lead time doubles?

c) How does your answer change if demand during lead time drops in half?
Answer #1

Given :Lead time = 21 days

Average daily usage = 100 units per day.

Appropriate reorder point = (average daily usage * lead time) + safty stock

assume safty stock = 0 as it is fast moving product

reorder point = 100 * 21 + 0

= 2100

When the inventory level falls to approximately 2100 units, a new order is to be placed. Notice that the reorder point is not related to the optimal order quantity or any of the inventory costs.

if demand during lead time doubles

If demand doubles ie 200 units per day

then reorder point = 200* 21 days

= 4200 units

When the inventory level falls to approximately 4200 units, a new order is to be placed.

If demand during lead time drops in half

then order point = 50*21

= 1050 units

When the inventory level falls to approximately 1050 units, a new order is to be placed.